Suppose that X, Y, and Z are jointly distributed random variables, that is, they are defined on the same sample space. Suppose that we also have the following. E(x)-5 ECY)4 E(Z)--8 Var (x)-39 Var (Y)-11 Var (z) 37 Compute the values of the expressions below. E(2 -2z) 35 30 Var (sz)-5-D
Answer #1

(a)

E(2-2Z) 2-2E(Z) = 2-(2x (-8)) = 2-16-18

So,

Answer is:

18

(b)

E( 3-3)

=rac{5}{3}E(Y)-rac{5}{3}E(Z)

20 40 20 (-1)--(-8) 6.6667

So,

Answer is:

6.6667

(c)

Var(SZ) _ 5 = 52Var(Z)-5 (25x 37) _ 5 925-5 920

So,

Answer is:

920

(d)

E(4Y^{2})=4E(Y^{2}) (1)

By Theorem:

Var(y) = E(y2 )-(E(y))

Substituting values, we get:

11=E(Y^{2})-(-4)^{2}=E(Y^{2})-16

So,

E(Y2) = 27

Substituting,(4) becomes:

EAY2)-16 × 27-432

So,

Answer is:

432